# Fundamental physical constants set the observability and operation of phase and other transitions and increase entropy

### v4

@@ Abstract
Approaching the problem of understanding fundamental physical constants (FPCs) started with discussing the role these constants play in high-energy nuclear physics and astrophysics. Condensed matter physics was relatively unexplored in this regard. More recently, it was realised that FPCs set lower or upper bounds on key condensed matter properties. In this Perspective, we discuss a wider role played by FPCs in condensed matter physics: at given environmental conditions, FPCs set the observability and operation of entire physical effects and phenomena. We discuss second-order phase transitions including structural and superconducting transitions and first-order transitions including melting and boiling. We also discuss metastable states, transitions between them, chemical reactions and their products. An interesting byproduct of this discussion is that the order of magnitude of the transition temperature can be calculated from FPCs. We show that the new states emerging as a result of various transitions increase the phase space and entropy. Were FPCs to take different values, these transitions would become inoperative at our environmental conditions and the new states due to these transitions would not emerge. This suggests that the current values of FPCs, by enabling various transitions and reactions which give rise to new states, promote entropy increase. We propose that this is a general effect that applies to different levels of structure in the Universe.
@@ 1 Introduction
Figure 1: Different scales of structure in the physical world including the range of typical sizes studied in condensed matter physics. Fairly recently, fundamental physical constants were shown to put lower and upper bounds on different condensed matter properties - properties that are already observed - including viscosity, diffusion and spin dynamics, electron and heat transport, thermal conductivity in insulators and conductors, the speed of sound, elastic moduli and velocity gradients that can be set up in cells using biochemical energy (see discussion in the Introduction). In section 3, we discuss a wider role of fundamental constants affecting observability and operation of entire phenomena and effects, including second-order structural and superconducting transitions, first-order melting transition, metastable transitions and chemical and other reactions.
@@
Understanding the values and origin of FPCs is one of the grandest challenges in modern science [14]. The problem of understanding fundamental constants is not that it involves very hard technical calculations but that we do not know anything more fundamental [15]. We do not know what kind of theoretical tools we need to address the problem. It is probably this enormity of the problem that explains why there seems to be no sustained programme of research in this area, scheduled meetings and so on, as compared to more developed fields of science.
@@
The top in Figure 1 involves cosmological scales and a balance between the gravitational constant and the cosmological constant. This balance gives enough time for the Universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form [6, 13].
For the rest of length scales in Figure 1, our theories imply that there is an intricate network of inter-related constraints on FPCs that need to be satisfied to give many structures in the observable Universe. This includes formation and evolution of different astronomical objects and in particular stars, synthesis of heavy elements in stars including carbon and oxygen and other elements which can then form molecular structures essential to life [6, 13]. The strength of these constraints, the degree of tuning between FPCs, depends on how well developed our physical models are. Ref. [13] discusses existing uncertainties involved in our models and how these uncertainties affect constraints imposed on FPC. It is important to note that FPCs involved in these processes are often the same and include the Planck constant, speed of light in vacuum, electron mass and charge and the proton-to-electron mass ratio (this is different from the constants operating at the very top and bottom scales mentioned earlier - these play role at their length and energy scales only). This means that the parameter space of these FPCs can be better constrained and understood.
@@
It was realised that FPCs set upper or lower bounds on several key condensed matter properties. Examples include bounds on viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26], thermal conductivity in insulators [27] and conductors [28], the speed of sound [29], elastic moduli including those in lower-dimensional systems [28, 30], velocity gradients that can be set up in cells using biochemical energy [31] and superconducting temperature [32]. This has deepened our understanding of these properties and their physical models.
In this Perspective, we discuss a wider role of FPCs: extending beyond setting bounds on properties that are already observed, fundamental constants affect observability and operation of entire physical phenomena and effects. In the next section, we discuss why discussing FPCs at the condensed matter scale is important. This is followed by considering second-order phase transitions including structural and superconducting transitions and first-order transitions including melting and boiling. This discussion shows how FPCs set the observability and operation of phase transitions and hence the existence of different states emerging as a result of these transitions, including metastable states. An interesting byproduct of this discussion is that the order of magnitude of the transition temperature in, for example, structural phase transitions, can be calculated from FPCs.
We show that the new states emerging from various transitions increase the phase space and entropy. We also show that were FPCs to take different values, these transitions would become inoperative at our environmental conditions, and the new statess would not emerge. This suggests that the current values of FPCs, by enabling various transitions and reactions which produce the new states, result in the increase of the phase space and entropy. We propose that this is a general effect that applies to other levels of structure in the Universe. This discussion invites an analogy with an evolutionary process where entropy increase selects for FPCs.
@@ 2 Why filling the condensed matter gap is important
There are several reasons why understanding the role of FPCs at condensed matter length and energy scales and filling the condensed matter gap in Figure 1 is important. First and quite generally, having a non-interrupted and continuous set of length and energy scales in Figure 1 over which we can discuss the role of FPCs gives more information and detail to use.
Second, recall that it is the uncertainty and insufficient development of physical models related to higher-energy processes that limits our understanding of FPCs [13]. On the other hand, condensed matter physics involves a larger variety of ways to test condensed matter theories. Many of these tests are more readily available than those in high energy physics. As a result, theories and concepts in condensed matter physics have been well verified and cross-checked using a great variety of experiments and simulations.
@@
The fourth and final reason discussed in this Perspective is that if we better understand the operation of FPCs in condensed matter physics, we may find that the underlying principles are quite general and can be applied to other length and energy scales. This can lead to new insights related to the origin of FPCs and their values.
In the next section, we discuss several examples where FPCs set observability and operation of several key condensed matter effects and phenomena.
3 Observability and operation of phase and other transitions. Emergence of new states and entropy
@@ 3.1 Fundamental and environmental parameters
We start with separating two types of parameters. The first type are FPCs. The second type are the environmental parameters such as temperature $T$ and pressure $P$ at typical planetary conditions where we do our observations. These parameters are not fundamental but are due to historical accidents [33] such as the distance between the planet and its star which affects the planet’s temperature. In the Solar system, for example, $T$ is in the range 102-103 K. In this section, we will consider how FPCs affect the observability and operation of condensed matter phenomena such as phase transitions and other reactions at our environmental conditions.
We will be discussing the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$ and the proton mass $m_{p}$. We will also be discussing two dimensionless ratios, the proton-to-electron mass ratio $\beta$
@@
$\alpha=\frac{e^{2}}{\hbar c}$ (2)
@@
$a_{\rm B}\approx 0.5$ and is on the shorter side of interatomic separations in condensed phases in the typical range 1-3 Å. $E_{\rm R}\approx$ 13.6 eV and is on the larger side of cohesive energies in condensed phases in the range 1-10 eV.
3.2 Second-order phase transitions
3.2.1 Enropy increase due to a phase transition
Our first case study are second-order phase structural transitions in solids. Let’s consider common silica SiO2 and its low-temperature phase, quartz. Silica has several polymorphs at different pressure and temperature, separated by second-order phase transitions. As a model example, let’s the high-temperature phase of silica, $\beta$-cristobalite, emerging at the phase transition [34, 35].
@@ 3.2.1 Enropy increase due to a phase transition
where $\bar{\omega}$ is the average phonon frequency defined as $\bar{\omega}^{3N}=\prod\limits_{i}\omega_{i}$, $\omega_{i}$ are $3N$ phonon frequencies and $N$ is the number of atoms. Here and below, $k_{\rm B}=1$.
@@
The entropy increase due to the phase transition can take place both above and below $T_{c}$. For reversible phase transition where the low-temperature phase is always recovered below $T_{c}$, the entropy increase takes place above $T_{c}$. In our model quartz-cristobalite example, this gives
@@
We find both quartz and cristobalite at our environmental conditions, hence let’s consider the entropy of the quartz-cristobalite mixture at any temperature below $T_{c}$ where both phases co-exist, $S_{Q-C}$. This mixture consists of any proportion of the two phases with the total number of atoms $N$. We compare $S_{Q-C}$ to the entropy of the initial quartz phase, $S_{Q}$, with the same number of atoms $N$. $S_{Q-C}$ is larger than $S_{Q}$ because the entropy of cristobalite is larger as discussed earlier. Hence,
@@
The increased phase space and entropy above $T_{c}$ become “frozen-in” in the metastable state below $T_{c}$.
@@
3.2.2 Observability of phase transitions. Entropy without a phase transition
@@
Let’s consider what would happen to a phase transition were FPCs to take different values. Let’s consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs. (12)-(13). Note that this question is well-posed because it is possible to vary $\omega_{m}$ and $\omega_{m}^{u}$ by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta$ and $\alpha$ in Eqs. (1) and (2) unchanged and hence maintain all preceding essential processes (e.g. stellar formation and evolution, nuclear synthesis, stability of ordered structures and so on) intact [6, 13].
Large $\omega_{m}$ implies large $\omega_{i}$ and $\omega_{\rm D}$ (recall that $\omega_{m}$ is close to $\omega_{\rm D}$) and $T\ll\hbar\omega_{\rm D}$ in the temperature range considered. This, in turn, implies that the high-temperature classical result (5) does not apply because the system is now in the quantum regime. In this regime, instead of Eq. (5), the entropy is [36]
@@
$S_{s}$ is close to 0 when $T\ll\hbar\omega_{\rm D}$. Then the second term in the free energy $F=E-TS$ is close to 0 too, and $F$ now does not have a crossover corresponding to a phase transition. The phase transition becomes inoperative.
@@
We see that the order of magnitude of $T_{c}$ (15) is set by FPCs. Were different values of FPCs to give $\omega_{m}$ and $T_{c}$ in Eq. (15) on the order of, for example, $10^{5}$ K or higher, phase transitions would be unobservable at our environmental conditions. This is consistent with our earlier result: large $\omega_{\rm D}$ gives small entropy in Eq. (14) and no phase transition.
From this and previous sections, we conclude that the current values of FPCs, by enabling the operation of phase transitions which give rise to new states, increase the phase space and entropy.
3.2.3 Summary
The discussion in Section 3.2 gives us several insights. First, FPCs govern key effects in condensed matter physics and define the regime in which the system is, classical or quantum. Second, FPCs affect the observability and operation of entire effects such as phase transitions. Third, FPCs, by enabling phase transitions, increase the phase space and entropy.
@@ 3.3 Superconducting transition
Using Eq. (17), we ask how would the observability of superconductivity change were fundamental constants to take different values? Since $T_{c}^{u}$ is the upper bound, $T_{c}^{u}$ on the order of 10-5 K or lower due to different FPCs implies that superconductivity would be unobserved.
@@
3.4 First-order phase transitions: solids, liquids, gases
@@
Figure 3: Three states of matter, solids, liquids and gases, separated by three phase transition lines. The triple and critical points are shown.
@@
We see that, similarly to second-order phase transitions discussed in Section 3.2, the current values of FPCs set the observability of first-order phase transitions in Figure 3 and the existence of different states of matter.
The apparent overlap of our environmental conditions and melting and boiling temperatures set by FPCs gives rise to two states important for life. First, H2O can exist as liquid water at our typical planetary conditions. Second, carbon-based molecules and compounds can exist as cohesive (bonded) structures. Much larger cohesive energy due to different values of FPCs would result in H2O existing as a solid only and hence no water-based life. Much smaller values of cohesive energies due to different FPCs would result in H2O and carbon-based structures not existing as cohesive states but as gas states only.
@@
Similarly to what we saw in Section 3.2 for the second-order transitions, the first-order phase transitions and the emerging high-temperature gas phase increase the entropy in comparison to what the entropy of the solid would have been in the absence of phase transitions (for first-order transitions, the entropy increase has a contribution from the latent heat). Similarly to the second-order transitions, the first-order transitions are enabled by FPCs as discussed earlier in this section.
Differently from solid and gas entropies we considered, discussing liquid entropy is more challenging theoretically due to the fundamental no-small parameter problem of the liquid theory (see Ref. [42] for a quick review and Refs. [45, 46] for a more detailed discussion). Common models such as the hard-sphere and Van der Waals models and their extensions give the constant-volume specific heat $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically rather than real liquids where $c_{v}=3$ close to melting [45, 46]. However, the energy, the primary property in statistical physics [36], can be calculated in liquids in the closed form using the main premise of statistical physics: the system’s energy is given by the energies of its excitations [47]. In liquids, these excitations are gapped phonons and particle jumps [45, 46, 48, 49, 50, 51]. The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}$ is calculated from the energy. Nevertheless, liquids are a condensed state of matter, and our discussion of how the interplay of cohesive energies set by FPCs and environmental parameters gives rise to different states of matter applies to liquids too.

### v5

# Fundamental physical constants, operation of physical phenomena and entropy increase

@@ Abstract
Approaching the problem of understanding fundamental physical constants (FPCs) started with discussing the role these constants play in high-energy nuclear physics and astrophysics. Condensed matter physics was relatively unexplored in this regard. More recently, it was realised that FPCs set lower or upper bounds on key condensed matter properties. Here, we discuss a much wider role played by FPCs in condensed matter physics: at given environmental conditions, FPCs set the observability and operation of entire physical effects and phenomena. We discuss structural and superconducting phase transitions and transitions between different states of matter, with implications for life processes. We also discuss metastable states, transitions between them, chemical reactions and their products. A byproduct of this discussion is that the order of magnitude of the transition temperature can be calculated from FPCs only. We show that the new states emerging as a result of various transitions increase the phase space and entropy. Were FPCs to take different values, these transitions would become inoperative at our environmental conditions and the new states due to these transitions would not emerge. This suggests that the current values of FPCs, by enabling various transitions and reactions which give rise to new states, promote entropy increase. Based on this entropy increase and the associated increase of statistical probability, we conjecture that entropy increase is a selection principle for FPCs considered to be variable in earlier discussions.
@@ 1 Introduction
Figure 1: Different scales of structure in the Universe including the typical size range in condensed matter physics. Fairly recently, FPCs were shown to put lower and upper bounds on different condensed matter properties - properties that are already observed - including viscosity, diffusion and spin dynamics, electron and heat transport, thermal conductivity in insulators and conductors, the speed of sound, elastic moduli, flow in cells, vibration frequency in solids and superconducting temperature (see Introduction). Here, we discuss a wider role of FPCs affecting observability and operation of entire phenomena and effects, including structural and superconducting transitions, transitions between different states of matter and chemical reactions.
@@
Understanding the values and origin of FPCs is one of the grandest challenges in modern science [14]. The problem of understanding FPCs is not that it involves very hard technical calculations but that we do not know anything more fundamental [15]. We do not know what kind of theoretical tools we need to address the problem. It is probably this enormity of the problem that explains why there seems to be no sustained programme of research in this area, scheduled meetings and so on, as compared to more developed fields of science.
@@
The top in Figure 1 involves cosmological scales and a balance between the gravitational and cosmological constants. This balance gives enough time for the Universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form [6, 13].
For the rest of intermediate length scales in Figure 1, our theories imply that there is an intricate network of inter-related constraints on FPCs that need to be satisfied to give many structures in the observable Universe. This includes formation and evolution of different astronomical objects and in particular stars, synthesis of heavy elements in stars including carbon and oxygen and other elements which can then form molecular structures essential to life [6, 13]. The strength of these constraints, the degree of tuning between FPCs, depends on how well developed our physical models are. Ref. [13] discusses existing uncertainties involved in our models and how these uncertainties affect constraints imposed on FPC. FPCs involved in different processes at intermediate scales are often the same and include the Planck constant, speed of light in vacuum, electron mass and charge and the proton-to-electron mass ratio (this is different from FPCs operating at the very top and bottom scales mentioned earlier - these play role at their length and energy scales only). This means that the parameter space of these FPCs can be better constrained and understood.
@@
It was realised that FPCs set upper or lower bounds on several key condensed matter properties. Examples include bounds on viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26, 27], thermal conductivity in insulators [28] and conductors [29], the speed of sound [30], elastic moduli including those in low-dimensional systems [29, 31], flow in cells [32, 33], vibration frequency in solids and superconducting temperature [34]. This has deepened our understanding of these properties and their physical models.
Here, we discuss a much wider role of FPCs in condensed matter physics: extending beyond setting bounds on properties that are already observed, FPCs affect observability and operation of entire physical phenomena and effects. In the next section, we discuss why discussing FPCs at the condensed matter scale is important. This is followed by considering phase transitions including structural and superconducting transitions and transitions between different states of matter with implications for life processes. This shows how FPCs set the operation of entire physical phenomena such as phase and other transitions and hence the existence of different states of matter emerging as a result of these transitions. An interesting byproduct of this discussion is that the order of magnitude of the transition temperature can be calculated from FPCs only.
We show that the new states emerging from various transitions increase the phase space and entropy. We also show that were FPCs to take different values, these transitions would become inoperative at our environmental conditions, and the new states would not emerge. This suggests that the current values of FPCs, by enabling various transitions and reactions which produce the new states, result in the increase of the phase space and entropy. We propose that this is a general effect operating at different levels of structure in the Universe.
The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where this increase is analogous to selection in biological evolution. Based on the entropy increase and the associated increase of statistical probability, we conjecture that the entropy increase is a selection principle for FPCs considered to be variable in earlier discussion [35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 33].
@@ 2 Why filling the condensed matter gap is important
There are several reasons why understanding the role of FPCs at condensed matter physics scales and filling the condensed matter gap in Figure 1 is important. First and quite generally, having a non-interrupted and continuous set of length and energy scales in Figure 1 over which we can discuss the role of FPCs gives more information and detail to use.
Second, the uncertainty and insufficient development of physical models related to higher-energy processes limits our understanding of FPCs [13]. On the other hand, condensed matter physics involves a larger variety of ways to test condensed matter theories. Many of these tests are more readily available than those in high energy physics. As a result, theories and concepts in condensed matter physics have been well verified and cross-checked using a great variety of experiments and simulations.
@@
The fourth and final reason discussed here is that if we better understand the operation of FPCs in condensed matter physics, we may find that the underlying principles are quite general and can be applied to other length and energy scales. This can lead to new insights related to the origin of FPCs and their values.
In the next sections, we discuss several examples where FPCs set observability and operation of several key condensed matter effects and phenomena.
3 Observability and operation of phase transitions. Emergence of new states with higher entropy
@@ 3.1 Fundamental and environmental parameters
[…]
We will be discussing FPCs central to condensed matter physics: the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$ and the proton mass $m_{p}$. As discussed in this section, the first three constants set the characteristic scales of distance and energy in condensed matter. $m_{p}$ sets the scale of lattice oscillation frequency (see section 3.2.2). We will also be discussing two dimensionless ratios, the proton-to-electron mass ratio $\beta$
@@
$\alpha=\frac{1}{4\pi\epsilon_{0}}\frac{e^{2}}{\hbar c}$ (2)
@@
$a_{\rm B}\approx 0.5$ and is on the shorter side of interatomic separations in condensed phases in the typical range 1–3 Å. $E_{\rm R}\approx$ 13.6 eV and is on the larger side of cohesive energies in condensed phases in the range 1–10 eV.
3.2 Structural phase transitions
3.2.1 Entropy increase due to a phase transition
Our first case study are structural phase transitions in solids. Let’s consider common silica SiO2 which has several polymorphs at different pressure and temperature, separated by phase transitions. As a model example, let’s consider two silica phases: quartz and high-temperature $\beta$-cristobalite phase [44, 45].
@@ 3.2.1 Entropy increase due to a phase transition
where $\bar{\omega}$ is the average phonon frequency set as $\bar{\omega}^{3N}=\prod\limits_{i}\omega_{i}$, $\omega_{i}$ are $3N$ phonon frequencies, $N$ is the number of atoms and anharmonic corrections are assumed small and not considered below. Here and below, $k_{\rm B}=1$.
@@
We consider $T_{c}$ to be within the range of environmental temperature, $T_{1}<T_{c}<T_{2}$. The entropy increase due to phase transitions can take place both above and below $T_{c}$. For reversible phase transition where the low-temperature phase is always recovered below $T_{c}$, the entropy increase takes place above $T_{c}$. In our model quartz-cristobalite example, this gives
@@
We find both quartz and cristobalite at our environmental conditions, hence let’s consider the entropy of the quartz-cristobalite mixture at any temperature below $T_{c}$ where both phases co-exist, $S_{Q-C}$. This mixture consists of any proportion of the two phases with the total number of atoms $N$. We compare $S_{Q-C}$ to the entropy of the initial quartz phase, $S_{Q}$, with the same number of atoms $N$. As discussed earlier, the entropy of cristobalite is higher than that of quartz. Hence,
@@
[…]
@@
3.2.2 Observability of phase transitions. Entropy changes.
@@
Let’s consider what would happen to a phase transition were FPCs to take different values. Let’s consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs. (12)-(13). Note that this question is well-posed because it is possible to vary $\omega_{m}^{u}$ by varying $m_{e}$ and $m_{p}$ or $\hbar$ and $e$ while keeping $\beta$ and $\alpha$ in Eqs. (1) and (2) unchanged and hence maintain all preceding essential processes (e.g. stellar formation and evolution, nuclear synthesis, stability of ordered structures and so on) intact [6, 13].
[…]
$T_{2}\ll\hbar\omega_{m}$ (14)
Recalling that $\omega_{m}$ is close to $\omega_{\rm D}$, condition (14) implies $T\ll\hbar\omega_{\rm D}$. This, in turn, implies that the high-temperature classical result (5) does not apply because the system is now in the quantum regime. In this regime, instead of Eq. (5), the entropy is [46]
@@
$S_{s}$ is close to 0 when $T\ll\hbar\omega_{\rm D}$. Then the second term in the free energy $F=E-TS$ is close to 0 too, and $F$ now does not have a crossover corresponding to a phase transition. The phase transition becomes inoperative, and there is no associated entropy increase discussed in the previous section.
@@
We see that the order of magnitude of $T_{c}$ (16) is set by FPCs. Were different values of FPCs to give $\omega_{m}$ and $T_{c}$ in Eq. (16) on the order of, for example, $10^{5}$ K or higher, phase transitions would be unobservable at our environmental conditions. This is consistent with our earlier result: large $\omega_{\rm D}$ gives small entropy in Eq. (15), no phase transition and no associated entropy increase.
We therefore find that FPCs, by enabling the operation of phase transitions which give rise to new states, increase the phase space and entropy. We note that phase transitions in the phase-changing subsystem (PCS) involve the heat flow from another, heat-supplying subsystem (HSS) in order to reach $T_{c}$ and the transition (if the second high-entropy phase in the PCS is metastable below $T_{c}$, the heat can be reversibly returned from the PCS to HSS below $T_{c}$, retaining the entropy increase in the metastable phase and involving little or no net heat exchange between PCS and HSS). The PCS and HSS are open, however the combined PCS and HSS subsystem can be considered closed. The entropy increase due to phase transitions in the PCS applies to this closed subsystem and represents the entropy increase which is extra as compared to the case where no new phases emerge in the PCS and the entropy increase is due to the trivial heat exchange between the PCS and HSS only.
In summary, the discussion in Section 3.2 has so far given us several insights:
  1. 1.
    FPCs govern key effects in condensed matter physics and define the regime in which the system is, classical or quantum.
  2. 2.
    FPCs set the observability and operation of entire physical phenomena such as phase transitions (more examples are in Sections 3.3, 3.4 and 3.5).
  3. 3.
    FPCs, by enabling phase transitions, increase the phase space and entropy (more discussion is in Sections 3.4, 3.5 and 4 below).
@@ 3.3 Superconducting transition
Using Eq. (18), we ask how would the observability of superconductivity change were fundamental constants to take different values? Since $T_{c}^{u}$ is the upper bound, $T_{c}^{u}$ on the order of 10-5 K or lower due to different FPCs implies that superconductivity would be unobserved at our experimental conditions.
@@
3.4 Transitions between states of matter: solids, liquids and gases
@@
Figure 3: Three states of matter (solids, liquids and gases) separated by three phase transition lines. Triple and critical points are shown.
@@
We see that, similarly to structural phase transitions discussed in Section 3.2, the current values of FPCs set the observability of transitions between different states of matter in Figure 3 and the existence of these states.
The apparent overlap of our environmental conditions and melting and boiling temperatures set by FPCs gives rise to two states important for life:
  1. 1.
    H2O can exist as liquid water at our typical planetary conditions.
  2. 2.
    Carbon-based molecules and compounds can exist as cohesive (bonded) structures.
Much larger cohesive energy due to different values of FPCs would result in H2O existing as a solid only and hence no water-based life. Much smaller values of cohesive energies due to different FPCs would result in H2O and carbon-based structures not existing as cohesive states but as gas states only.
A more detailed consideration shows an additional and more specific role of FPCs in life processes: FPCs have a biofriendly window constrained by biofriendly viscosity and diffusion setting the motion in essential processes in and across cells [32, 33].
@@
Similarly to what we saw in Section 3.2 for structural phase transitions, the emerging high-temperature gas phase increases the entropy in comparison to what the entropy of the solid would have been in the absence of phase transitions (for first-order transitions, the entropy increase has a contribution from the latent heat). Similarly to structural transitions, transitions between different states of matter are enabled by FPCs as discussed earlier in this section.
Differently from solid and gas entropies in Eqs. (5), (15) and (20), no equation for the liquid entropy exists due to the fundamental no-small parameter problem of the liquid theory (see Ref. [52] for a quick review and Refs. [55, 56] for a more detailed discussion). Common models such as the hard-sphere and Van der Waals models and their extensions give the constant-volume specific heat $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically rather than real liquids where $c_{v}=3$ close to melting [55, 56]. However, the energy, the primary property in statistical physics [46], can be calculated in liquids in the closed form using the main premise of statistical physics: the system’s energy $E$ is set by the energies of its excitations [57]. In liquids, these excitations are gapped phonons and particle jumps [55, 56, 58, 59, 60, 61]. The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}=\frac{dE}{dT}$ (in almost entire range of liquid relaxation time, $E$ and $c_{v}$ are vibrational to a very good approximation [62]). Nevertheless, liquids are a cohesive state, and our discussion of how different states of matter emerge depending on cohesive energies set by FPCs applies to liquids too.